Are there any values of p such that p2+48 is equal to -14p?


1
Expert's answer
2022-06-01T07:58:03-0400

If p2+48 is equal to -14p then we have the next equation:

"p^2+48=-14p,"

"p^2+14p+48=0."

To solve this quadratic equation, let's find its discrininant:

"D=b^2-4ac=14^2-4\cdot 1\cdot 48=196-192=4."

As "D>0" the equation has two real distinct roots:

"p_1=\frac{-b-\sqrt{D}}{2a}=\frac{-14-2}{2}=-8,"

"p_2=\frac{-b+\sqrt{D}}{2a}=\frac{-14+2}{2}=-6."

So, there are two values of p: -8 and -6, such that p2+48 is equal to -14p:

At "p=-8:"

"(-8)^2+48=64+48=112=-14\cdot (-8)."

At "p=-6:"

"(-6)^2+48=36+48=84=-14\cdot (-6)."


Answer: p=-8 and p=-6.